OpenStudy (anonymous):

Write the expression in standard form. 3/(3-12i)

OpenStudy (jdoe0001):

what would be the conjugate of the denominator?

OpenStudy (anonymous):

3+12i

OpenStudy (jdoe0001):

yes, so let's use that, to multiply the rational by 1, a/b * 1 = a/b, so is fine, so let's do that :)

OpenStudy (jdoe0001):

$$\bf \color{blue}{\cfrac{3+12i}{3+12i} =1}\\ \cfrac{3}{3-12i} \times \cfrac{3+12i}{3+12i}\\ \cfrac{3(3+12i)}{(3-12i)(3+12i)}\\ \text{keep in mind that}\\ (a-b)(a+b) = (a^2-b^2)\\ $$ so what would that give you in the denominator?

OpenStudy (anonymous):

9-(12i^2) which is 9+144

OpenStudy (jdoe0001):

yes

OpenStudy (jdoe0001):

so that would give you \(\bf \cfrac{9+36i}{153} \implies \cfrac{9}{153}+\cfrac{36i}{153}\)

OpenStudy (jdoe0001):

so that would be the so-called standard form of "ax+bi"

OpenStudy (anonymous):

thanks